OpenStudy (anonymous):

-2sin(x/2) write in a asinb(x-h) form

zepdrix (zepdrix):

So I guess you're writing it in that form so you can identify the pieces, yes? \[\large -2\sin\left(\frac{x}{2}\right) \qquad = \qquad \color{orangered}{-2}\sin\left(\color{royalblue}{\frac{1}{2}}x\right)\] \[\large \color{orangered}{a}\sin\left(\color{royalblue}{b}x\right)\] Can you identify the \(\large a\) and \(\large b\) and \(\large h\) ? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you please help me/thank you!!

zepdrix (zepdrix):

I rewrote \(\large \dfrac{x}{2}\) as \(\large \dfrac{1}{2}x\). They reason we're allowed to do this is because there is always an invisible 1 on top. \[\large \frac{x}{2} \qquad = \qquad \frac{1\cdot x}{2} \qquad = \qquad \frac{1}{2}x\] Are you confused about the a and b...? I color-coded them. I thought that would have made it easier to identify them.

OpenStudy (anonymous):

what i mean how to find the value of h?

OpenStudy (anonymous):

please help me

zepdrix (zepdrix):

\[\large \color{orangered}{-2}\sin\left(\color{royalblue}{\frac{1}{2}}x\right) \qquad =\qquad \color{orangered}{-2}\sin\left[\color{royalblue}{\frac{1}{2}}(x+0)\right]\]

zepdrix (zepdrix):

Ok now it's in the form that they want. What is your h value?

OpenStudy (anonymous):

0?

zepdrix (zepdrix):

yesssss

OpenStudy (anonymous):

thank you